OpenStudy (anonymous):

The total number of students and adults at a concert was 460. The cost of tickets was $12 per student and $16 per adult. The total amount of income from tickets was $6520. How many students and how many adults were at the concert?

OpenStudy (anonymous):

let the no of students = x then amount collected from these x students = 12 * x = 12x Let the number of adults = y Then amount collected from these y adults = 16 * y = 16y From the above we get two equations x + y = 460 ----------(1) 12x + 16y = 6520 ----------(2) Rearranging (1) we get x = 460-y ------------------(3) Substituting this in (2) we get 12(460-y) + 16y = 6520 5520 - 12y + 16y = 6520 4y = 6520 - 5520 y = 1000/4 y = 250 putting this value of y in (3) we get x= 460 - 250 = 210 so number of students was = x = 210 number of adults was = y = 250